Dimensions are in meter

2. EVALUATE THE INDICATED BAR FORCES IN THE PIN CONNECTED TRUSSES : 3OAN kuwa SOAN 40kN + 4 d. 4 J 20kN TORN 10kN
Answer #1

(EM) = 0 e = tant (3) x coso(s) + 10(4)+20(9)=0 *(4) 6+ 70+ 160=0 20KN lokn LOKN X = - 41.6 KN Am 4m = 41.6 KN (compression)we have to solve one question per answer. please upload the other one seperately.